suppose you have a certain land area to be use for planting vegetable. you have 800 m of fencing materials and you want to use it for rectangular piece of land with an area of 30 000 square m, what will be the dimension of the fenced land?

Answer by richwmiller(17219)

2l+2w=800
l+w=400,
l*w=30000
l = 100, w = 300
or
l = 300, w = 100